OpenStudy (anonymous):

Use the Zero Product Property to solve the equation: -3m(2m – 5) = 0

jagr2713 (jagr2713):

is there and answer choice

jimthompson5910 (jim_thompson5910):

Hint: If A*B = 0 then A = 0 or B = 0

OpenStudy (anonymous):

Is it 225? I have no clue..

jimthompson5910 (jim_thompson5910):

if -3m(2m – 5) = 0 then -3m = 0 or 2m - 5 = 0

jimthompson5910 (jim_thompson5910):

so you need to solve these equations -3m = 0 or 2m - 5 = 0

OpenStudy (anonymous):

Am I solving for M?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

And the answer wouldn't be zero? For the second equation i got 5/2..

jimthompson5910 (jim_thompson5910):

either m = 0 or m = 5/2

jimthompson5910 (jim_thompson5910):

those are the two solutions